The New England Merchants Bank Building in Boston is 152 m high. On windy days it sways with a frequency of 0.15Hz , and the acceleration of the top of the building can reach 2.5% of the free-fall acceleration, enough to cause discomfort for occupants.

 

What is the total distance, side to side, that the top of the building moves during such an oscillation?


Answer is 2A=55 cm. How do we get to this answer?
Answer #1

a = - w^2 x

0.025*9.81 = (2*pi*0.15)^2*x

x=0.276 m

so 2x = 0.55 m = 55 cm